OpenStudy (anonymous):

Write an equation of a line with the given slope and y-intercept. m = –5, b = –3

OpenStudy (anonymous):

y = -5x - 3

OpenStudy (anonymous):

A. y = –5x – 3 B. y = –5x + 3 C. y = 5x – 3 D. y = –3x – 5

OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1)

OpenStudy (anonymous):

A. 3/2 B. -2/3 C. -3/2 D. 2/3